WestGas Conveyance, Inc., is a large U.S. natural gas pipeline company that wants to raise $120 million to finance expansion. WestGas wants a capital structure that is 50% debt and 50% equity. Its corporate combined federal and state income tax rate is 40%. WestGas finds that it can finance in the domestic U.S. capital market at the rates listed below. Both debt and equity would have to be sold in multiples of $20 million, and these cost figures show the component costs, each, of debt and equity if raised 50% by debt and 50% by equity. A London bank advises WestGas that U.S. dollars could be raised in Europe at the following costs, also in multiples of $20 million, while maintaining the 50/50 capital structure. Each increment of cost would be influenced by the total amount of capital raised. That is, if WestGas first borrowed $20 million in the European market at 5% and matched this with an additional $20 million of equity, additional debt beyond this amount would cost 11% in the United States and 10% in Europe. The same relationship holds for equity financing.

a. Calculate the lowest average cost of capital for each increment of $40 million of new capital, where WestGas raises $20 million in the equity market and an additional $20 in the debt market at the same time.

b. If WestGas plans an expansion of only $60 million, how should that expansion be financed? What will be the weighted average cost of capital for the expansion?

Rates:

Cost of

Domestic

Equity

Cost of

Domestic

Debt

Cost of

European

Equity

Cost of

European

Debt

Up to​ $40 million of new capital

12​%

7​%

13​%

5​%

​ $41 million to​ $80 million of new capital

17​%

11​%

15​%

10​%

Above​ $80 million

23​%

15​%

24​%

16​%
Answer #1
Tax rate 40%
Cost of Cost of After tax Cost of Cost of Cost of After tax Cost of
Domestic Domestic Domestic European European European
Equity Debt Debt Equity Debt Debt
Up to​ $40 million of new capital 12% 7% 4.20% 13% 5% 3.00%
​ $41 million to​ $80 million of new capital 17% 11% 6.60% 15% 10% 6.00%
Above​ $80 million 23% 15% 9.00% 24% 16% 9.60%
Lowest costs
Up to​ $40 million of new capital Domestic equity + European debt =0.5*12%+0.5*3% 7.50%
​ $41 million to​ $80 million of new capital European equity + European debt =0.5*15%+0.5*6% 10.50%
Above​ $80 million Domestic equity + Domestic debt =0.5*23%+0.5*9% 16.00%

Assume that the debt to equity ratio of 1:1 needs to maintained for the new capital fo $60M. Up to $40M, European debt+ US equity is the cheapest. From $41M to $60M, European debt and European equity is the cheapest. Hence, cost structure is as below

Funding source Value ($M) Cost
European Debt 20 3%
US Equity 20 12%
European Debt 10 6%
European Equity 10 15%
Total 60 8.50%